OpenStudy (anonymous):

Write a rule for the linear function in the table. x f(x) 1 3 2 2 3 1 4 0

OpenStudy (anonymous):

can you rewrite this? i'm confused

OpenStudy (anonymous):

it want let me put the chart

OpenStudy (anonymous):

A. f(x) = 4x B. f(x) = x – 4 C. f(x) = x + 4 D. f(x) = 4 – x

OpenStudy (anonymous):

so what's your real question? just find a rule? or find the values of f(x) by plugging the values of x?